# CORE - A Cell-Level Coarse-to-Fine Image Registration Engine for Multi-stain Image Alignment

### v1

@@ 1 Introduction
In computational pathology, extensive research has been directed towards biomarker discovery, characterisation of tissue architecture, and disease sub-typing through the analysis of multi-stained whole slide images (WSI). Since raw tissue samples are inherently transparent, they are typically stained with haematoxylin and eosin (H&E), immunohistochemistry (IHC), multiplex immunofluorescence (mIF), or periodic acid-Schiff (PAS) after fixation on glass slides. These stained samples are then scanned to generate high-resolution digital WSIs, which are then used for downstream analyses, such as biomarker identification and investigation of disease progression. To achieve a comprehensive understanding of tissue morphology and pathology, it is often necessary to integrate information from multiple stains. This requires precise WSI registration to enable joint analysis across different staining modalities. Image registration refers to the process of determining a transformation that maps corresponding elements of a source image to a target image. This transformation maps both images originating from different coordinate systems into a common coordinate space allowing for an accurate and meaningful comparison.
@@
Different stains and imaging modalities such as H&E, IHC, PAS and mIF highlight distinctive tissue components, leading to variations in contrast and texture. As a result, features that are prominent in one stain (or modality) may be faint or absent in another, making feature matching extremely challenging. Additionally, tissue sections often undergo both elastic deformations (such as stretching) and inelastic deformations (such as tear or missing regions) during preparation, further complicating alignment. Variability in section thickness and tissue handling can also change the appearance of the same tissue across slides, adding another layer of complexity. The digitisation process may introduce artefacts including blurriness, dust, and folds that degrade image quality and can mislead registration algorithms. Moreover, the high resolution of WSIs demands substantial computational power and storage, which poses practical challenges for efficient processing. Unlike conventional medical images such as MRI or CT scans, WSIs typically lack consistent landmarks across stains and modalities, making classic computer vision image registration techniques less effective. These challenges necessitate the need for specialised algorithms capable of robust and efficient registration, even in the presence of significant appearance and structural variability. Our work addresses these limitations by proposing a stain-agnostic, nuclei-point based registration framework that generalises across staining protocols without manual tuning. We formulate registration as a two-stage process: a coarse alignment based on prompt-guided tissue segmentation masks, followed by a fine-grained nuclei-level point set registration. This hybrid approach enables robust alignment across stain combinations (e.g., H&E–mIF, H&E-IHC, H&E-PAS) while preserving tissue structure and achieving high accuracy.
@@
High resolution visualisation - We provide real-time, high-resolution visualisation of both rigid and non-rigid deformations across the full WSIs, facilitating intuitive assessment and validation. A working demo of CORE registration is available at TiaViz Demo111https://tiademos.dcs.warwick.ac.uk/bokeh_app?demo=WSIReg
Open-source code for reproducibility - We provide open-source code through GitHub 222https://github.com/eshasadia/WSI_mIF_Registration.git to support reproducibility and future research and development in multimodal WSI registration.
@@ 2 Related Works
During the past decade, WSI registration has attracted significant research attention, leading to noteworthy contributions [lotz_comparison_2023, wodzinski_deeperhistreg_2024, gatenbee_virtual_2023]. Despite these advances, the unique characteristics of multi-stained WSIs, including their gigapixel resolution, tissue variability, non-linear deformations, and staining inconsistencies, continue to pose major challenges. These factors often limit the effectiveness of conventional medical image registration methods, particularly when alignment at the nuclei level is required. A detailed review of the WSI registration methods has been presented in Elhaminia et al. [elhaminia_traditional_2025].
@@
Notably, several works have explored registration through features derived from segmented tissues. For example, Ge et al. [ge_unsupervised_2022] developed a multi-stain registration method that employed an unsupervised structural feature-guided (SFG) CNN, which was robust to both low-resolution rough and high-resolution fine structural features of tissues. The non-rigid network was pre-trained with the synthetic Flying Chair dataset [dosovitskiy2015flownet]. Then the supervised component of the SFG network was trained on landmarks. Mahapatra et al. [mahapatra_registration_2020] leveraged deep segmentation maps from a pre-trained U-Net to guide non-linear alignment. These approaches highlight the trend towards integrating tissue morphology, nuclei features, and learning-based representations.
@@ 3 Methodology
We propose a two-stage framework for multimodal WSI registration, designed to progressively refine spatial alignment across multiple resolutions. The proposed approach comprises (i) a coarse registration and (ii) a fine shape-aware registration stage for low and high resolution WSI alignment.
@@ 3.1.1 Preprocessing
The preprocessing pipeline begins by converting RGB images to greyscale, followed by gamma correction to enhance the visibility of poorly stained tissue features. For each image pair, downsampled thumbnails are extracted at and resolutions. Differential gamma correction is employed to enhance tissue contrast, with gamma values of – for H&E-stained images and – for IHC-stained images. These ranges were empirically determined based on evaluations conducted across the datasets. Stain normalisation is then applied to improve tissue mask extraction in prompt-based segmentation.
@@
To achieve consistent colour representation across WSIs, we apply Macenko stain normalisation [macenko_method_2009], a well-established approach for standardising the appearance of H&E WSIs.
We follow the original parameter settings proposed by Macenko (, , ) and apply the reference H&E stain vectors, H: [0.650, 0.704, 0.286] and E: [0.072, 0.990, 0.105] to ensure robust and consistent stain separation and normalisation. The preprocessing pipeline incorporates normalisation techniques to overcome inter-sample variability caused by inconsistent staining and imaging conditions. This approach is essential for producing standardised images, allowing the model to focus on relevant structural features without being influenced by colour discrepancies supporting accurate segmentation outcomes.
@@ 3.1.3 Non-Rigid Registration
To address complex non-rigid spatial deformations present in WSIs, we propose a hierarchical deformation-based registration framework. Our framework builds on established hierarchical strategies used in both classical approaches (e.g., ANTs [avants_reproducible_2011]) and learning-based models (e.g., VoxelMorph [balakrishnan_voxelmorph_2018], DeeperHistReg [wodzinski_deeperhistreg_2024]), which estimate deformations in a coarse-to-fine manner to ensure stable convergence and improved capture of multi-scale variation.
@@
The registration algorithm is implemented using grid-based spatial transformations and bilinear interpolation. At each pyramid level, the displacement field is refined and propagated to the next higher resolution via upsampling to 6 levels in the pyramid. The modular design of the framework allows fine-grained control over resolution depth, optimiser parameters, and regularisation strength at each level. Our experiments across a diverse range of tissue sections and staining modalities confirm the robustness of the proposed approach. The method reliably aligns WSIs affected by physical deformation, staining variability, and scanning inconsistencies.
For deformation estimation, we employ NCC as the primary similarity metric for cost estimation during each iteration:
@@ 3.2 Fine-Grained Shape-Aware Registration
For local-level deformation estimation, we propose a fine-grained, shape-aware registration framework designed to address tissue-level deformations by aligning nuclei centroids extracted from WSIs. The method aims to enable precise correspondence at the nuclear level between source and target nuclei point sets, for both consecutive and re-stained tissue sections. Fig. 3 shows fine-grained shape-aware pointset registration block for nuclei level WSI registration.
@@
Next, H-maxima and H-minima transforms are used to detect local intensity extrema corresponding to nuclear regions. These extrema serve as seed points for a marker-controlled watershed segmentation. The watershed algorithm operates on the distance transform of the binary mask, using the detected minima as markers, enabling effective separation of clustered or overlapping nuclei while preserving precise boundary definitions.
@@
Shape-aware Alignment: This step further refines coarsely aligned WSI pair for global transformations, including translation, rotation, and scaling using nuclei points.
@@ 3.2.1 Shape-aware Alignment
where is the shape weight parameter, is the transformation, and , are normalised shape attributes.
@@
We use Powell’s derivative-free optimisation. Iteration continues until:
where is the small convergence threshold, and is the maximum iteration count in Equation 8. Table A.3 lists all the parameters used for shape-aware alignment along with it’s range and deafult value.
@@
: Achieves an optimal balance between morphology and other factors.
@@
When working with re-stained slides, we performed shape-aware alignment using the complete set of detected nuclei points at the highest resolution. This approach is effective because the underlying tissue architecture remains physically identical across staining iterations, allowing for direct point-to-point correspondence matching with minimal spatial variation.
@@ 3.2.2 Local Deformation Estimation
To provide CPD with reliable initial correspondences, we employ a mutual nearest neighbours (MNN) approach that enforces bidirectional consistency. This removes ambiguous matches and enhances the robustness of the registration. MNNs are computed by identifying pairs of points that are closest to each other between the two sets. Specifically, the algorithm first finds the nearest point in the moving set for each fixed point, and then the nearest point in the fixed set for each moving point. A pair is considered a mutual nearest neighbour if each point is the closest to the other. Table A.4 reports all the parameters used for non-rigid registration along with range and default values.
@@ 4.1.3 HYRECO- Hybrid Re-stained and Consecutive Histological Serial Sections
The HYRECO dataset comprises two complementary subsets. Subset A (consecutive section slides) contains 9 groups of consecutive tissue sections stained with H&E, CD8 (Cluster of Differentiation 8), CD45 (Cluster of Differentiation 45), and Ki67. PHH3 (Phosphohistone H3) stained slides were generated by bleaching and re-staining the original H&E sections using a t-CyCIF-like technique. Each section includes 11–19 manually verified landmarks, resulting in 138 landmarks per stain and 690 in total. Subset B (re-stained slides) consists of 54 paired H&E–PHH3 images, containing a total of 2,303 annotated landmarks (approximately 43 per pair).
@@ 4.1.4 Multi-IHC CRC
The Multi-IHC CRC dataset, obtained from the University Hospitals Coventry and Warwickshire (UHCW) NHS Trust, comprising WSIs from 86 distinct patients [awan_deep_2023]. Each case features 16 high-resolution slides scanned at 0.275 microns/pixel with various stains including CK8/18 (Cytokeratin 8 and Cytokeratin 18), mismatch repair (MMR) markers (MLH1, MSH2, MSH6, PMS2), and H&E. For evaluation of registration methods, we selected 7 cases with 6 slides per case for direct comparison with Awan et al. 2023 [awan_deep_2023]. A notable challenge is the significant structural variability between H&E and IHC slides (separated by approximately 50 m), while consecutive IHC sections maintain higher structural correlation.
@@ 4.1.5 REACTIVAS
The REACTIVAS dataset was acquired from the University Hospitals Birmingham NHS Foundation Trust. It consists of Renal biopsy samples stained using a variety of protocols, including H&E, PAS, and advanced mIF techniques. We have included 11 paired samples for all 3 staining protocols. The mIF modality offers rich cellular and molecular detail, enabling the simultaneous visualisation of multiple biomarkers within the same tissue section. This complements the structural information provided by traditional histological stains, enhancing the dataset’s value for integrative pathology research.
@@
| REACTIVAS | Renal biopsies, H&E, PAS, mIF | 11 paired samples | 0.2 |
@@ 4.3 Results
Table 2 presents a direct comparison between the proposed CORE method and current state-of-the-art approaches on the ACROBAT dataset [weitz_acrobat_2024], demonstrating superior performance and reduced computational time. The proposed method attains higher accuracy, as indicated by the AMTRE (90th percentile) metric, while requiring significantly less processing time than competing techniques. Fig. 4 illustrates representative source (IHC) and target (H&E) slides from the ACROBAT dataset, together with the extracted masks, registration results before and after alignment, and the corresponding extracted patch and full-resolution image.
@@
Table 3 presents a comparison between the proposed CORE method and existing state-of-the-art models on the ANHIR dataset [borovec_anhir_2020], evaluated at both the coarse and fine registration stages. The results demonstrate improved performance and reduced runtime at coarse level. Fine-shape aware registration was performed only on slides available at 20 and 40 magnifications, while lower-resolution samples (1.25, 2.5, 5, and 10) were excluded fine resolution registration. Table A.5 reports step-wise results on the ANHIR dataset, summarising performance before initial alignment, after coarse alignment, and following fine alignment. Fig. 5 illustrates representative source (ER) and target (H&E) slides from the ANHIR dataset, along with the extracted masks, registration results before and after alignment, and the corresponding extracted patch and full-resolution image.
@@
Table 4 present rigid and non-rigid comparative results of state-of-the-art models and the proposed CORE method on the HYRECO dataset, encompassing both re-stained and consecutive sections. In terms of median TRE for rigid registration, the proposed method performs comparably to existing state-of-the-art approaches, surpassing all except HistokatFusion [lotz_comparison_2023]. However, it achieves substantially lower runtime across all comparisons, providing a faster and more reliable registration. For non-rigid registration, the proposed method outperforms all competing approaches. Table A.6 summarises step-wise results on the HYRECO dataset, detailing performance before initial alignment, after coarse alignment, and following fine alignment. Fig. 6 illustrates registration results on slides from the HYRECO dataset, showing comparisons before and after alignment. Fig. 11 presents landmarks overlaid on the HYRECO source and target images. From left to right, the panels depict the source image, the target image, and the landmarks overlaid on the target before initial alignment, after coarse registration, and following fine registration. Fig. 7 displays registration results on a representative patch extracted from the HYRECO dataset. From left to right, the panels show the overlay before registration, after coarse registration, after fine-rigid registration, and after fine non-rigid registration. Figure A.2 shows registration overlay on HYRECO patches extracted at full resolution.
@@
Following evaluation on public datasets, the proposed CORE method was further tested on private datasets. Table 5 presents a comparison between CORE and publicly available state-of-the-art models on the Multi-IHC CRC dataset, demonstrating superior performance at both coarse and fine registration stages, as indicated by lower TRE and reduced runtime. Table A.7 summarises step-wise results on the Multi-IHC CRC dataset, detailing performance before initial alignment, after coarse alignment, and following fine alignment. Fig. 8 illustrates visual registration results on Multi-IHC CRC slides before and after alignment. Fig. 10 shows landmarks overlaid on Multi-IHC CRC source and target images. From left to right, the panels depict the source, the target, and the landmarks overlaid on the target before initial alignment, after coarse registration, and after fine registration.
@@
| AMTRE [m] | AATRE [m] | Avg Time (s) | AMTRE [m] | Avg Time (s) | |||
| CORE | 65.26 | 109.83 | 18 | 6.81 | 18 | ||
| Fine | CORE | 39.20 | 60.21 | 2340 | 2.26 | 2340 | |
| NonRigid | HistokatFusion [lotz_comparison_2023] | 5.30 | - | 1831 | 0.90 | 1831 | |
| DeeperHistReg [wodzinski_regwsi_2024] | 4.96 | – | – | 0.59 | – | ||
| CORE | 4.35 | – | 2850 | 0.41 | 2850 | ||

### v3

@@ Abstract
Keywords:
Registration, Microscopy, Multimodal, Nuclei, Alignment, Morphology, Deformation
@@ 1 Introduction
In computational pathology, extensive research has been directed towards biomarker discovery, characterisation of tissue architecture, and disease sub-typing through the analysis of multi-stained whole slide images (WSI). Since raw tissue samples are inherently transparent, they are typically stained with haematoxylin and eosin (H&E), immunohistochemistry (IHC), multiplex immunofluorescence (mIF), or periodic acid-schiff (PAS) after fixation on glass slides. These stained samples are then scanned to generate high-resolution digital WSIs, which are then used for downstream analyses, such as biomarker identification and investigation of disease progression. To achieve a comprehensive understanding of tissue morphology and pathology, it is often necessary to integrate information from multiple stains. This requires precise WSI registration to enable joint analysis across different staining modalities. Image registration refers to the process of determining a transformation that maps corresponding elements of a source image to a target image. This transformation maps both images originating from different coordinate systems into a common coordinate space, allowing for an accurate and meaningful comparison.
@@
Different stains and imaging modalities such as H&E, IHC, PAS and mIF highlight distinctive tissue components, leading to variations in contrast and texture. As a result, features that are prominent in one stain (or modality) may be faint or absent in another, making feature matching extremely challenging. Additionally, tissue sections often undergo both elastic deformations (such as stretching) and inelastic deformations (such as tear or missing regions) during preparation, further complicating alignment. Variability in section thickness and tissue handling can also change the appearance of the same tissue across slides, adding another layer of complexity. The digitisation process may introduce artefacts, such as blurriness, dust, and folds that degrade image quality and can mislead registration algorithms. Moreover, the high resolution of WSIs demands substantial computational power and storage, which poses practical challenges for efficient processing. Unlike conventional medical images such as MRI or CT scans, WSIs typically lack consistent landmarks across stains and modalities, making classic computer vision image registration techniques less effective. These challenges necessitate the need for specialised algorithms capable of robust and efficient registration, even in the presence of significant appearance and structural variability. Our work addresses these limitations by proposing a stain-agnostic, nuclei-point based registration framework that generalises across staining protocols without manual tuning. We formulate registration as a two-stage process: a coarse alignment based on prompt-guided tissue segmentation masks, followed by a fine-grained nuclei-level point set registration. This hybrid approach enables robust alignment across stain combinations (e.g., H&E–mIF, H&E-IHC, H&E-PAS) while preserving tissue structure and achieving high accuracy.
@@
High resolution visualisation - We provide real-time, high-resolution visualisation of both rigid and non-rigid deformations across the full WSIs, facilitating intuitive assessment and validation. A working demo of CORE registration is available at TiaViz Demo111https://tiademos.dcs.warwick.ac.uk/bokeh_app?demo=WSIReg. Details for using TiaViz have been provided in Appendix A
Open-source code for reproducibility - We provide open-source code through GitHub 222https://github.com/eshasadia/WSI_mIF_Reg.git to support reproducibility and future research and development in multimodal WSI registration.
@@ 2 Related Works
During the past decade, WSIs registration has attracted significant research attention, leading to noteworthy contributions. Despite these advances, the unique characteristics of multi-stained WSIs, including their gigapixel resolution, tissue variability, non-linear deformations, and staining inconsistencies, continue to pose major challenges. These factors often limit the effectiveness of conventional medical image registration methods, particularly when alignment at the nuclei level is required. A detailed review of the WSI registration methods has been presented in Elhaminia et al. [elhaminia_traditional_2025].
@@
Notably, several works have explored registration through features derived from segmented tissues. For example, Ge et al. [ge_unsupervised_2022] developed a multi-stain registration method that employed an unsupervised structural feature-guided (SFG) CNN, which was robust to both low-resolution rough and high-resolution fine structural features of tissues. The non-rigid network was pre-trained with the synthetic Flying Chair dataset [dosovitskiy2015flownet]. Then the supervised component of the SFG network was trained on landmarks. Mahapatra et al. [mahapatra_registration_2020] leveraged deep segmentation maps from a pre-trained U-Net to guide non-linear alignment.
These approaches highlight the trend towards integrating tissue morphology, nuclei features, and learning-based representations.
@@ 3 Methodology
We propose a two-stage framework for multimodal WSI registration, designed to progressively refine spatial alignment across multiple resolutions. The proposed approach comprises (i) a coarse registration and (ii) a fine shape-aware registration stage for high resolution WSI alignment.
@@ 3.1.1 Preprocessing
The preprocessing pipeline begins by converting RGB images to greyscale, followed by gamma correction to enhance the visibility of poorly stained tissue features. For each WSI pair, downsampled thumbnails are extracted at and resolutions. Differential gamma correction is employed to enhance tissue contrast, with gamma values of – for H&E-stained images and – for IHC-stained images. These ranges were empirically determined based on evaluations conducted across the datasets.
Stain normalisation is then applied to improve tissue mask extraction for prompt-based segmentation.
@@
To achieve consistent colour representation across WSIs, we apply Macenko stain normalisation [macenko_method_2009], a well-established approach for standardising the appearance of H&E WSIs. We follow the original parameter settings proposed by Macenko (, , ) and apply the reference H&E stain vectors, H: [0.650, 0.704, 0.286] and E: [0.072, 0.990, 0.105] to ensure robust and consistent stain separation and normalisation. The preprocessing pipeline incorporates normalisation techniques to overcome inter-sample variability caused by inconsistent staining and imaging conditions. This approach is essential for producing standardised images, allowing the model to focus on relevant structural features without being influenced by colour discrepancies supporting accurate segmentation outcomes. However, the model does not rely solely on stain normalization. Tissue mask extraction can still perform effectively without stain normalisation, provided that the tissue remains clearly visible, a condition guaranteed by the preceding contrast enhancement stage. Further details are provided in Appendix A.
@@ 3.1.3 Non-Rigid Registration
To address complex non-rigid spatial deformations present in WSIs, we propose a hierarchical deformation-based registration framework. Our framework builds on established hierarchical strategies used in both classical approaches (e.g., ANTs [avants_reproducible_2011]) and learning-based models (e.g., VoxelMorph [balakrishnan_voxelmorph_2018], DeeperHistReg [wodzinski_deeperhistreg_2024]), which estimate deformations in a multi-resolution manner to ensure stable convergence and improved capture of multi-scale variations.
@@
The registration algorithm is implemented using grid-based spatial transformations and bilinear interpolation. At each pyramid level, the displacement field is refined and propagated to the next higher resolution via upsampling to 6 levels in the pyramid up to 1.25 resolution. The modular design of the framework allows precise control over resolution depth, optimiser parameters, and regularisation strength at each level. Our experiments across a diverse range of tissue sections and staining modalities confirm the robustness of the proposed approach. The method reliably aligns WSIs affected by physical deformation, staining variability, and scanning inconsistencies.
For deformation estimation, we employ local NCC as the primary similarity metric with window size of 7 for cost estimation during each iteration:
@@ 3.2 Fine-Grained Shape-Aware Registration
For local-level deformation estimation, we propose a fine-grained, shape-aware registration framework designed to address tissue-level deformations by aligning nuclei centroids extracted from WSIs at 20 or 40. The method aims to enable precise correspondence at the nuclear level between source and target nuclei point sets, for both consecutive and re-stained tissue sections. Fig. 3 shows fine-grained shape-aware pointset registration block for nuclei level WSI registration.
@@
Next, H-maxima and H-minima (morphological operations used to filter local maxima and minima) transforms are used to detect local intensity extrema corresponding to nuclear regions. These extrema serve as seed points for a marker-controlled watershed segmentation. The watershed algorithm operates on the distance transform of the binary mask, using the detected minima as markers, enabling effective separation of clustered or overlapping nuclei while preserving precise boundary definitions.
@@
Shape-aware Alignment: This step further refines coarsely aligned WSI pairs for global transformations, including translation, rotation, and scaling using nuclei points.
@@ 3.2.1 Shape-aware Alignment
where is the shape weight parameter, is the transformation, and , are normalised shape attributes estimated from shape area i.e per nuclei area corresponding to the coordinates.
@@
We use Powell’s derivative-free optimisation and iteration continues until:
where is the small convergence threshold, and is the maximum iteration count in Equation 8. Table A.3 lists all the parameters used for shape-aware alignment along with it’s range and default values.
@@
: Achieves an optimal balance between morphology and extreme transformation.
@@
When working with re-stained slides, we performed shape-aware alignment using the complete set of detected nuclei points at the highest resolution. This approach is effective because the underlying tissue architecture remains physically identical across staining iterations, allowing for direct nuclei-to-nuclei correspondence matching with minimal spatial variation.
@@ 3.2.2 Local Deformation Estimation
To provide CPD with reliable initial correspondences, we employ a mutual nearest neighbours (MNN) approach that enforces bidirectional consistency. This removes ambiguous matches and enhances the robustness of the registration. MNNs are computed by identifying pairs of points that are closest to each other between the two sets. Specifically, the algorithm first finds the nearest point in the source set for each target point, and then the nearest point in the target set for each source point. A pair is considered a mutual nearest neighbour if each point is the closest to the other. Table A.4 reports all the parameters used for non-rigid registration along with range and default values.
@@ 4.1.3 HYRECO- Hybrid Re-stained and Consecutive Histological Serial Sections
The HYRECO dataset comprises two complementary subsets. Subset A (consecutive section slides) contains 9 groups of consecutive tissue sections stained with H&E, CD8 (Cluster of Differentiation 8), CD45 (Cluster of Differentiation 45), and Ki67. PHH3 (Phosphohistone H3) stained slides were generated by bleaching and re-staining the original H&E sections using a t-CyCIF (tissue-based Cyclic Immunofluorescence) like technique. Each section includes 11–19 manually verified landmarks, resulting in 138 landmarks per stain and 690 in total. Subset B (re-stained slides) consists of 54 paired H&E–PHH3 images, containing a total of 2,303 annotated landmarks (approximately 43 per pair).
@@ 4.1.4 Multi-IHC CRC
The Multi-IHC CRC dataset, obtained from the University Hospitals Coventry and Warwickshire (UHCW) NHS Trust, comprising WSIs from 86 distinct patients [awan_deep_2023]. Each case features 16 high-resolution slides scanned at 0.275 µm/px with various stains including CK8/18 (Cytokeratin 8 and Cytokeratin 18), mismatch repair (MMR) markers (MLH1, MSH2, MSH6, PMS2), and H&E. For evaluation of registration methods, we selected 7 cases with 6 slides per case for direct comparison with Awan et al. 2023 [awan_deep_2023]. A notable challenge is the significant structural variability between H&E and IHC slides (separated by approximately 50 m), while consecutive IHC sections maintain higher structural correlation.
@@ 4.1.5 REACTIVAS
The REACTIVAS dataset [web:reactivas] was acquired from the University Hospitals Birmingham NHS Foundation Trust. It consists of Renal biopsy samples stained using a variety of protocols, including H&E, PAS, and advanced mIF techniques. We have included 11 paired samples for all 3 staining protocols. The mIF modality offers rich cellular and molecular detail, enabling the simultaneous visualisation of multiple biomarkers within the same tissue section. This complements the structural information provided by traditional histological stains, enhancing the dataset’s value for integrative pathology research.
@@
| REACTIVAS | Renal biopsies, H&E, PAS, mIF | 11 paired samples | 0.28 |
@@ 4.3 Results
Table 2 presents a direct comparison between the proposed CORE method and current state-of-the-art approaches on the ACROBAT dataset [weitz_acrobat_2024], demonstrating superior performance and reduced computational time. The proposed method attains higher accuracy, as indicated by the AMTRE (90th percentile) metric, while requiring significantly less processing time than competing techniques. Fig. 4 illustrates representative target (H&E), source (IHC) and registered source slides from the ACROBAT dataset, together with the registration results overlay before and after alignment, and the corresponding extracted patch at full-resolution.
@@
Table 3 presents a comparison between the proposed CORE method and existing state-of-the-art models on the ANHIR dataset [borovec_anhir_2020], evaluated at both the coarse and fine registration stages. The results demonstrate improved performance and reduced runtime at coarse level. Fine-shape aware registration was performed only on slides available at 20 and 40 magnifications, while lower-resolution samples (1.25, 2.5, 5, and 10) were excluded fine resolution registration. Table A.6 reports step-wise results on the ANHIR dataset, summarising performance before initial alignment, after coarse alignment, and following fine alignment. Fig. 5 illustrates representative target (H&E), source (ER) and registered slides from the ANHIR dataset, along with the registration results overlay before and after alignment, and the corresponding extracted patch at full-resolution.
@@
Table 4 presents comparative results of state-of-the-art models and the proposed CORE method on the HYRECO dataset, encompassing both re-stained and consecutive sections. In terms of the final median TRE for registration, the proposed method outperforms existing state-of-the-art approaches. We have also presented coarse and fine level rigid registration results of CORE and comparison with state-of-the-art approaches in Table A.5. It achieves substantially lower runtime across all comparisons at coarse level, providing a faster and more reliable registration. Table A.7 summarises step-wise results on the HYRECO dataset, detailing performance before initial alignment, after coarse alignment, and following fine alignment. Fig. 6 illustrates representative target (H&E), source (PHH3) and registered slides from the HYRECO dataset, along with the registration results overlay before and after alignment, and the corresponding extracted patch at full-resolution.
@@
Following evaluation on public datasets, the proposed CORE method was further tested on private datasets. Table 5 presents a comparison between CORE and publicly available state-of-the-art models on the Multi-IHC CRC dataset, demonstrating superior performance at both coarse and fine registration stages, as indicated by lower TRE and reduced runtime. Table A.8 summarises step-wise results on the Multi-IHC CRC dataset, detailing performance before initial alignment, after coarse alignment, and following fine alignment. Fig. 7 illustrates representative target (CK818), source (MLH1) and registered slides from the Multi-IHC CRC dataset, along with the registration results overlay before and after alignment, and the corresponding extracted patch at full-resolution.
@@
| Resolution | Technique | Consecutive Sections | Re-stained Sections | |||
| AMTRE [m] | AATRE [m] | Avg Time (s) | AMTRE [m] | Avg Time (s) | ||
| Fine | HistokatFusion [lotz_comparison_2023] | 5.30 | - | 1831 | 0.90 | 1831 |
| DeeperHistReg [wodzinski_regwsi_2024] | 4.96 | – | – | 0.59 | – | |
| CORE | 4.35 | – | 2850 | 0.41 | 2850 | |